OpenStudy (anonymous):

Linda's paycheck was for $500.00. When she cashed her check at the bank, the teller gave her a total of 30 bills, both tens and twenties. How many twenties did the teller give Linda?

OpenStudy (ash2326):

Let the teller gave her x tens and y twenties so There were totally 30 bills, so \[x+y=30\] We also know that amount was $500, so \[10\times x+20\times y=500\] Now we have two equations and two variables, can you solve this ?